OpenStudy (anonymous):

what is the simplified form of y^2-6y+8/ y^2-7y+12

OpenStudy (anonymous):

\[y-2 \over y-3\] because \[y^2-6t+8 = (y-4)(y-2)\] and\[ y^2-7y+12 = (y-4)(y-3)\] Cancel out the (y-4)s and that's your answer.